OpenStudy (anonymous):

Prove that X^(2n) - y^(2n) is divisible by (x+y)? Mathamatical Induction

OpenStudy (anonymous):

@apoorvk @RaphaelFilgueiras

OpenStudy (anonymous):

PLz read the question Now

terenzreignz (terenzreignz):

Let's factor it! :D \[x^{2n} - y^{2n} = (x^{n} + y^{n})(x^{n} - y^{n})\] \[x^{n} - y^{n} = (x - n)(x^{n-1} + x^{n-2}y ......xy^{n-2} + y^{n-1})\] Therefore: \[x^{2n} - y^{2n} = (x^{n} + y^{n})(x - y)(x^{n-1} + x^{n-2}y......xy^{n-2} + y^{n-1})\] And clearly \[(x - y)\] is a factor :) QED Cheers :D

terenzreignz (terenzreignz):

Oh cr*p, I misread 0.0 Not sure how to prove (x+y) is a factor.... I'll have to pass on that one T.T

terenzreignz (terenzreignz):

Only if n is odd

OpenStudy (callisto):

For n=1 x^[2(1)] - y^[2(1)] = x^2 - y^2 = (x+y)(x-y), which is divisible by x+y Therefore, the proposition is true for n=1. Assume that for some positive integers k, x^(2k) - y^(2k) is divisible by x+y, that is x^(2k) - y^(2k) = m (x+y) , where m is a positive integer. When n = k+1 x^[2(k+1)] - y^[2(k+1)] = x^2 [x^(2k)] - y^2 [y^(2k)] = x^2 [m (x+y) +y^(2k)] - y^2 [y^(2k)] = m(x+y) x^2 + x^2 y^(2k) - y^2 [y^(2k)] = m(x+y) x^2 + y^(2k) (x^2 - y^2) = m(x+y) x^2 + y^(2k) (x-y)(x+y) = (x+y) [ mx^2 + y^(2k) (x-y)] , where [ mx^2 + y^(2k) (x-y)] is a positive integer. Therefore x^[2(k+1)] - y^[2(k+1)] is divisible by (x+y) Therefore, the proposition is also true for n=k+1 By principle of mathematical induction, the proposition is true for all positive integers n.

terenzreignz (terenzreignz):

@Callisto One thing's for sure, I didn't deserve that medal @Yahoo! This is the real solution, mine was totally WRONG

OpenStudy (callisto):

@terenzreignz You can't hardly judge if you deserve the medal....